OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)? Q(1, 11) R(2, 4) S(4, -4) T(9, -2)

OpenStudy (muzzack):

WHATs the qeustion

OpenStudy (anonymous):

I posted it

OpenStudy (anonymous):

Which point lies on a circle with a radius of 5 units and center at P(6, 1)?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

The answer of this question is "R(2,4)".

OpenStudy (anonymous):

your an angel thank you

OpenStudy (anonymous):

The equation of this circle in this problem is \[(x - 6)^{2} + (y - 1)^{2} = 5^{2}\] When you substitute R(2,4), \[(2-6)^{2} + (4-1)^{2} = (-4)^{2} + 3^{2} = 16 + 9 = 25 = 5^{2}\] Do you understand it?